OpenStudy (anonymous):

suppose you randomly choose an integer n between 1 to 5, and then draw a circle with a radius of n centimetres. What is the expected area of this circle to the nearest hundredths of a square centimetres?

OpenStudy (anonymous):

area of the circle is \(\pi r^2\) so your expected values is \[\frac{1}{5}\times \left(\pi +4\pi+9\pi+16\pi+25\pi\right)\]

OpenStudy (anonymous):

how did you get those values? Can you explain further please.